# Online Resource Allocation via Static Bundle Pricing

### v1

@@ Appendix D Graph Setting
$\mathbb{E}_{M,I}\left[L_{(s,t)}(e)\right]\leq\left\lfloor x_{w}\right\rfloor\cdot\sum_{p\in\mathcal{P}_{s,t}^{e}}\frac{x_{w,p}}{x_{w}}+\sum_{p\in\mathcal{P}_{s,t}^{e}}x_{>w,p}\leq\sum_{b\in\mathcal{A}}\sum_{p\in\mathcal{P}_{s,t}^{e}}\sum_{v\in\mathcal{S}}x_{b,v,p}$
Summing over all desired $(s,t)$ paths concludes the proof. Moreover, in each case $L_{(s,t)}(e)$ is equal or upper bounded by a sum of independent or negatively associated (NA) $\{0,1\}$ random variables. Indeed, for Cases $1,3,4$ the proof is straightforward.
@@
$\Big\{\mathbbm{1}{\left\{\text{$w$-copy offered}\right\}}\cdot\mathbbm{1}{\left\{\exists b:\,v_{b}=w\right\}},\ \{\mathbbm{1}{\left\{v_{b}>w\right\}}\}_{b\in\mathcal{A}}\Big\}\ \text{is NA.}$
@@ D.2 Subadditivity
We now relax the simplifying assumption that a buyer who desires an $(s,t)$-path is only interested on $(s,t)$-path copies. The proof is similar to the single-minded case. Similarly, a buyer could purchase several paths $p_{1},\ldots,p_{\ell}$, that correspond to $(s_{1},t_{1}),\ldots,(s_{\ell},t_{\ell})$ types respectively, from the posted menu such that $\bigcup_{i}p_{i}\supseteq p\in\mathcal{P}_{s,t}$. We will show that after applying essentially the same simple normalization step to the fractional LP solution as in the single-minded case, no buyer can ever strictly prefer such a cover over purchasing an $(s,t)$-path directly. In other words, the posted prices are subadditive.
@@
$\textsc{FracWeight}(x)=\sum_{b}\sum_{v\in\mathcal{S}}\sum_{p\in\mathcal{P}_{b}}x_{b,v,p}\quad\text{and}\quad w_{e}(x)=\sum_{b}\sum_{v\in\mathcal{S}}\sum_{p\in\mathcal{P}_{b}\,:\,e\in p}x_{b,v,p}.$
We now state the equivalent of Lemma˜B.1 for the graph setting.
@@
Fix a desired $(s,t)$-path such that $x_{v}<q_{v}$. Suppose there exist different $(s_{i},t_{i})$-paths $p_{1},\dots,p_{\ell}$ such that $\bigcup_{i=1}^{\ell}p_{i}\supseteq p$, for some $p\in\mathcal{P}_{s,t}$. If $x_{(s_{i},t_{i}),v_{i},p_{i}}>0$ for all $i\in[\ell]$, then
@@
for some $p\in\mathcal{P}_{s,t}$. Note that since we always offer $|\mathcal{A}_{s,t}|$ copies of $(s,t)$-paths at some price, a copy of an $(s,t)$-path will always be available as in the single-minded setting.
First, if $r(S)=1$ then $r(T_{i})=0$, contradiction. Now notice that $x_{T_{i},r(T_{i})}>0$ and $x_{S,r(S)-1}$ is not tight by the definition of our pricing rules. Now note that we cannot directly apply Lemma˜D.1 as in the single-minded setting since $x_{T_{i},r(T_{i})}>0$ does not directly imply $x_{T_{i},r(T_{i}),p_{i}}>0$. However, we can overcome this issue via a mass-transfer argument.
Indeed, if $r(T_{i})$ is a crucial value for $T_{i}$ then since path $p_{i}$ is realized we get $x_{T_{i},r(T_{i}),p_{i}}>0$ from the definition of the roulette-drawing, thus we are done in this case. Suppose now that $r(T_{i})$ is tight and $x_{T_{i},r(T_{i}),p_{i}}=0$. Then, since $p_{i}$ is realized it holds that $x_{T_{i},\geq r(T_{i}),p_{i}}>0$ by the definition of the roulette drawing. Thus, there exists a value $v>r(T_{i})$ such that $x_{T_{i},v,p_{i}}>0$. Also, since $r(T_{i})$ is tight there exists a path $p_{i}^{\prime}$ such that $x_{T_{i},r(T_{i}),p_{i}^{\prime}}>0$. We now transfer sufficiently small mass from $x_{T_{i},v,p_{i}}>0$ to $x_{T_{i},r(T_{i}),p_{i}}=0$ and from $x_{T_{i},r(T_{i}),p_{i}^{\prime}}>0$ to $x_{T_{i},v_{i},p_{i}^{\prime}}$. Note that these transfers do not interfere with the feasibility, optimality or the preprocessing step, since they occur between same paths and we balance the value exchanges. Thus, we can assume that $x_{T_{i},r(T_{i}),p_{i}}>0$ for each $i\in[\ell]$.
As a result, we can apply Lemma˜D.1 and get contradiction as in the single-minded case.
@@ Appendix E Lower Bounds
Here we show how to remove the $\ln m$ loss in this special case by using Theorem 3 in [poljak]. The theorem states that if $t$ is a prime power and $t^{2}-t$ divides ntn-t, then
@@
We restrict to the case where the number of items is a cube of a prime, $m=n=p^{3}$ for a sufficiently large prime $p$, and set $t\coloneqq p,\ r\coloneqq 2$. Since $t=p$ is a prime (hence a prime power) and $t^{2}-t=p^{2}-p$ divides $n-t=p^{3}-p$, the condition of Theorem 3 in [poljak] holds, and we obtain
$g(n,t,2)\;\geq\;t^{\frac{n-t}{t^{2}-t}}\;=\;p^{p+1}=t^{t+1}.$
@@
We now reuse exactly the same construction and analysis as in the proof of Theorem 1.4, specialized to $B=1$ ($r=2$), with this choice of $m$ and $t$. Lemma G.3 gives $\mathbb{P}\!\left[X\geq 1\right]\;\geq\;0.5$, since $Np=B$.
The structural part is exactly as in the $B$-copies proof, with $B=1$ and $r=2$: the analogue of Observation 7.4 (for $B=1$) says that no feasible allocation can serve buyers from more than one group; the analogue of Observation 7.5 (with $B=1$) says that if $E_{i}$ holds, the offline optimum can serve all $t$ buyers in group $i$. Thus, repeating the offline argument from Theorem 1.4 with $B=1$,
@@
Similarly, the same argument from the proof of Theorem 1.4 with $B=1$ shows that each group has total expected value $1$ and, once the algorithm takes value from a group, the expected remaining value in that group is at most 1; combined with the fact that at most one group can be partially served when $B=1$, this yields
@@
which proves Theorem 1.6.
E.2 Proof of Theorem 1.7
We first note that one can obtain a lower bound of this form by translating the lower-bound instance of Saxena, Velusamy, and Weinberg [SVW22] for prophet inequalities over $q$ partition matroids into a $q$-single-minded instance and counting the number of items.
Instead of going through this translation, we give a direct combinatorial proof, by connecting the Alon–Alweiss measure (AAM) to qualitatively independent partitions and then plugging the resulting partitions into the single-minded lower-bound template used in the proof of Theorem 1.6.
@@
Let $\ell:=\textsc{AAM}(p,N)$. By definition there exists a $\mathbb{Z}_{p}$-covering $(\ell,p)$-family $\mathcal{C}\subseteq\mathbb{Z}_{p}^{\ell}$ with $|\mathcal{C}|\geq N$, i.e. for all distinct $u,v\in\mathcal{C}$ and all $s\in\mathbb{Z}_{p}$ there is $i\in[\ell]$ such that $u_{i}-v_{i}\equiv s\pmod{p}$. Set $X:=[\ell]\times\mathbb{Z}_{p}$, so $|X|=p\ell$. For each $u=(u_{1},\dots,u_{\ell})\in\mathcal{C}$ define a $p$-partition $\mathcal{P}_{u}$ of $X$ by blocks
@@
hence $(i,a+u_{i})=(i,b+v_{i})\in B_{v}(b)$ and so $B_{u}(a)\cap B_{v}(b)\neq\emptyset$. Since $a,b$ were arbitrary, $\mathcal{P}_{u}$ and $\mathcal{P}_{v}$ are qualitatively independent. Thus $\{\mathcal{P}_{u}:u\in\mathcal{C}\}$ is a family of $|\mathcal{C}|\geq N$ pairwise qualitatively independent $p$-partitions of an $n=p\ell=p\cdot\textsc{AAM}(p,N)$ element set. Therefore $g\bigl(p\cdot\textsc{AAM}(p,N),\,p,\,2\bigr)\geq N$. ∎
Saxena, Velusamy, and Weinberg [SVW22] show that there exists an absolute constant $\alpha>0$ such that for all sufficiently large primes $p$,
@@
We then instantiate exactly the single-minded lower-bound construction from the proof of Theorem 1.6 for $B=1$ using this family $\mathcal{F}$. The structural and probabilistic analysis of this instance (number of "fully active" groups, offline prophet value, and the bound on any online algorithm) is identical to the $B=1$ case of Theorem 1.6 and implies:
@@
Appendix F Tree Setting
In this section, we show how to achieve an $O(1)$ competitive ratio for the case where single-minded buyers demand paths in unit-capacitated trees and have $\{0,1\}$ valuations. While this result is already known due to Chawla et al. [chawla18] (see their more general Theorem 1.4), we include it here to illustrate how our framework can recover existing results, thereby demonstrating its versatility and power.
Formally, consider a tree $T=(V,E)$ with unit capacities on all edges and set an arbitrary node $r\in V$ to be its root. Each buyer’s value is realized according to a Bernoulli distribution, i.e., $\mathcal{D}_{b}=\mathrm{Be}(q_{b})$. Using the same techniques as in [chawla18], one could partition buyers’ values into $\log(v_{\max}/v_{\min}^{*})$ valuation classes to achieve their $O(\log H)$ competitive ratio. For clarity, however, we do not perform this step here, since it is straightforward and already established; we focus on the analysis as a proof of concept to illustrate our framework.
Approach.
The approach follows our general algorithmic framework and the analysis borrows the main ideas from [chawla18]. First, solve the ex-ante linear relaxation with appropriately scaled capacities (Step 1). For trees, we select $\gamma=10$. Next, post the randomized menu resulting from Step 2. We will now prove that the above (randomized) menu obtains a constant competitive ratio.
Definitions.
We say that a buyer $b$ is blocked by another buyer $b^{\prime}\neq b$ if $S_{b}$ and $S_{b^{\prime}}$ share an edge and $b^{\prime}$ is accepted by the constrained algorithm before $b$ arrives. Let $\textnormal{{Alg}}_{M,I}$ be the set of buyers accepted by the constrained algorithm, and define $\textsc{Blocked}_{M,I}\;=\;\textnormal{{UAlg}}_{M,I}\setminus\textnormal{{Alg}}_{M,I},$ the set of buyers accepted by the unconstrained algorithm but rejected by the constrained one due to blocking.
For a buyer $b$, let $u_{b}$ be the node of $S_{b}$ that is closest to the root $r$. Any edge of $S_{b}$ incident to $u_{b}$ is called a peak edge of $b$; the remaining edges of $S_{b}$ are non-peak. The parent of an edge is the unique adjacent edge closer to the root. Let $e_{b}^{1}$ and $e_{b}^{2}$ denote the peak edges of $b$ (one or two, depending on the structure of $S_{b}$), and define $f_{b}$ as the unique edge that is the parent of $e_{b}^{1}$; if $e_{b}^{2}$ also exists, then this edge is also the parent of $e_{b}^{2}$. We partition blocked buyers according to where the blocking occurs:
  • $\textsc{Peak}_{M,I}$ is the set of buyers in $\textsc{Blocked}_{M,I}$ that are blocked on one of their peak edges,
  • $\textsc{NonPeak}_{M,I}$ is the set of buyers in $\textsc{Blocked}_{M,I}$ that are blocked only on non-peak edges.
Finally, let $\ell_{A}(e)$ denote the load of edge $e$ due to buyers in $A$. Note that $\ell_{\textnormal{{UAlg}}_{M,I}}(e)=L_{M,I}(e)$.
Analysis.
Combining Lemma˜3.4 and Lemma˜3.3 with our choice of $\gamma=10$ (and using the fact that $c(e)=1$ for all edges), we obtain the following.
Corollary F.1.
For every edge $e\in E$,
$\mathbb{E}_{M,I}\left[L_{M,I}(e)\right]\;\leq\;\frac{c(e)}{10}\;=\;0.1.$
Moreover,
$\mathbb{E}_{M,I}\left[\textnormal{{UAlg}}_{M,I}\right]\;\geq\;\frac{\textnormal{{FracOpt}}_{\gamma}}{8}\;\geq\;\frac{\textnormal{{FracOpt}}}{80}.$
For a fixed $M$ and $I$ it holds that (note that buyers are $\{0,1\}$ thus only the cardinality of the sets matters):
${}\textnormal{{Alg}}_{M,I}=\textnormal{{UAlg}}_{M,I}-\textsc{Blocked}_{M,I}=\textnormal{{UAlg}}_{M,I}-\textsc{Peak}_{M,I}-\textsc{NonPeak}_{M,I}$ (2)
We first prove the following lemma.

### v2

@@ Appendix D Graph Setting
$\mathbb{E}_{M,I}\left[L_{(s,t)}(e)\right]\leq\left\lfloor x_{(s,t),w}\right\rfloor\cdot\sum_{p\in\mathcal{P}_{s,t}^{e}}\frac{x_{(s,t),\geq w,p}}{x_{(s,t),\geq w}}+q_{(s,t),>w}\sum_{p\in\mathcal{P}_{s,t}^{e}}\frac{x_{(s,t),\geq w,p}}{x_{(s,t),\geq w}}\leq\sum_{b\in\mathcal{N}_{s,t}}\sum_{p\in\mathcal{P}_{s,t}^{e}}\sum_{v\in\mathcal{U}}x_{b,v,p}$
Summing over all desired $(s,t)$ pairs concludes the proof. Moreover, in each case $L_{(s,t)}(e)$ is equal or upper bounded by a sum of independent or negatively associated (NA) $\{0,1\}$ random variables. Indeed, for Cases $1,3,4$ the proof is straightforward.
@@
$\Big\{\mathbbm{1}{\left\{\text{$w$-copy offered}\right\}}\cdot\mathbbm{1}{\left\{\exists b:\,v_{b}=w\right\}},\ \{\mathbbm{1}{\left\{v_{b}>w\right\}}\}_{b\in\mathcal{N}_{s,t}}\Big\}\ \text{is NA.}$
@@ D.2 Subadditivity
We now relax the simplifying assumption that a buyer who desires an $(s,t)$-pair is only interested in $(s,t)$-path copies. The proof is similar to the single-minded case. Similarly, a buyer could purchase several paths $p_{1},\ldots,p_{\ell}$, that correspond to $(s_{1},t_{1}),\ldots,(s_{\ell},t_{\ell})$ types respectively, from the posted menu such that $\bigcup_{i}p_{i}\supseteq p\in\mathcal{P}_{s,t}$. We will show that after applying essentially the same simple normalization step to the fractional LP solution as in the single-minded case, no buyer can ever strictly prefer such a cover over purchasing an $(s,t)$-path directly. In other words, the posted prices are subadditive.
@@
$\textsc{FracWeight}(x)=\sum_{b}\sum_{v\in\mathcal{U}}\sum_{p\in\mathcal{P}_{b}}x_{b,v,p}\quad\text{and}\quad w_{e}(x)=\sum_{b}\sum_{v\in\mathcal{U}}\sum_{p\in\mathcal{P}_{b}\,:\,e\in p}x_{b,v,p}.$
We now state the equivalent of Lemma B.1 for the graph setting.
@@
Fix a desired $(s,t)$-path such that $x_{(s,t),v}<q_{(s,t),v}$. Suppose there exist different $(s_{i},t_{i})$-paths $p_{1},\dots,p_{\ell}$ such that $\bigcup_{i=1}^{\ell}p_{i}\supseteq p$, for some $p\in\mathcal{P}_{s,t}$. If $x_{(s_{i},t_{i}),v_{i},p_{i}}>0$ for all $i\in[\ell]$, then
@@
for some $p\in\mathcal{P}_{s,t}$. Note that since we always offer $|\mathcal{N}_{s,t}|$ copies of $(s,t)$-paths at some price, a copy of an $(s,t)$-path will always be available as in the single-minded setting.
First, if $r(S)=1$ then $r(T_{i})=0$, contradiction. Now notice that $x_{T_{i},r(T_{i})}>0$ and $x_{S,r(S)-1}$ is not tight by the definition of our pricing rules. Now note that we cannot directly apply Lemma D.1 as in the single-minded setting since $x_{T_{i},r(T_{i})}>0$ does not directly imply $x_{T_{i},r(T_{i}),p_{i}}>0$. However, we can overcome this issue via a mass-transfer argument.
Suppose that $x_{T_{i},r(T_{i}),p_{i}}=0$. Then, since $p_{i}$ is realized it holds that $x_{T_{i},\geq w_{T_{i}},p_{i}}>0$ by the definition of the roulette drawing (recall that $w_{S}$ denotes the important value of $S$). Thus, there exists a value $v$ such that $x_{T_{i},v,p_{i}}>0$. Also, since $r(T_{i})$ is a price there exists a path $p_{i}^{\prime}$ such that $x_{T_{i},r(T_{i}),p_{i}^{\prime}}>0$. We now transfer sufficiently small mass from $x_{T_{i},v,p_{i}}>0$ to $x_{T_{i},r(T_{i}),p_{i}}=0$ and from $x_{T_{i},r(T_{i}),p_{i}^{\prime}}>0$ to $x_{T_{i},v,p_{i}^{\prime}}$. Note that these transfers do not interfere with the feasibility, optimality or the preprocessing step, since they occur between same paths and we balance the value exchanges. Thus, we can assume that $x_{T_{i},r(T_{i}),p_{i}}>0$ for each $i\in[\ell]$.
As a result, we can apply Lemma D.1 and get contradiction as in the single-minded case.
@@ Appendix E Lower Bounds
We give a variant of the lower-bound template from Theorem 2.4 that additionally enforces a maximum desired bundle size $d$. First we introduce the notion of balanced $r$-way QI partitions.
Balanced r-way QI Partitions.
We now refine Poljak’s probabilistic argument to the balanced setting. The only difference is that we sample partitions from uniformly random balanced labelings.
Lemma E.1 (Balanced refinement of Theorem 4 in Poljak and Tuza [1989]).
Let $n,t,r\geq 1$ be integers with $t\mid n$, and set $d:=n/t$. Then
$g_{\mathrm{bal}}(n,t,r,d)\;\geq\;\frac{r}{et}\cdot\exp\!\left(\frac{n}{rt^{r}}\right),$
Proof.
For each $u\in U$ and each $\ell\in[r]$,
$\mathbb{P}\!\left[u\in P^{(j_{\ell})}_{a_{\ell}}\right]=\frac{d}{n}=\frac{1}{t},$
and by independence across the $r$ sampled partitions,
Hence
$\mathbb{E}\left[Z\right]=\frac{n}{t^{r}}.$
Moreover, for a fixed partition $j$ and class $a$, $P^{(j)}_{a}$ is a uniformly random subset of $U$ of size $d$. It is a well known fact that for a uniformly random size-$d$ subset, the membership indicators are negatively associated (sampling without replacement) therefore
Similarly to Poljak and Tuza [1989], choosing
$k=\left\lceil\frac{r}{et}\cdot\exp\!\left(\frac{n}{rt^{r}}\right)\right\rceil$
We now set $r:=B+1$, since $(B+1)$-way QI is exactly what rules out serving buyers from more than $B$ groups when there are $B$ copies of each item (as in Claim 10.4). We will take $m=dt$ items and require that every class in our partitions has size at most $d$; with $t\mid m$, balanced sampling gives classes of size exactly $m/t=d$.
Define, for $d$ sufficiently large and $1\leq B<\ln d$,
Lemma E.2.
With $t$ and $m$ as in (5),
$g_{\mathrm{bal}}(m,t,B+1,d)\;\geq\;B\,t^{\,t},$
for all sufficiently large $d$.
Proof.
Apply Lemma E.1 with $n=m=dt$, $r=B+1$, and $d=n/t$:
$g_{\mathrm{bal}}(m,t,B+1,d)\;\geq\;\frac{(B+1)}{et}\cdot\exp\!\left(\frac{m}{(B+1)t^{B+1}}\right)=\frac{(B+1)}{et}\cdot\exp\!\left(\frac{d}{(B+1)t^{B}}\right).$
By the definition of $t$,
Hence
$g_{\mathrm{bal}}(m,t,B+1,d)\;\geq\;d^{\Omega(t)}.$
Lower-Bound Instance with Bundle Cap d.
Let $m=dt$ items and suppose there are exactly $B$ copies of each item. By Lemma E.2, there exists a $(B+1)$-way QI family
of $d$-balanced partitions of the $m$ items into $t$ classes each:
$\mathcal{P}^{(i)}=\{P^{(i)}_{1},\dots,P^{(i)}_{t}\},\qquad|P^{(i)}_{a}|\leq d.$
Analysis.
Define events $E_{i}$ and the random variable $X$ exactly as in Theorem 2.4. Then $X\sim\mathrm{Bin}(N,(1/t)^{t})$ and $Np=B$, so
$\mathbb{P}\!\left[X\geq B\right]\geq 0.5$
by the proof of Lemma F.4 (see Case 2).
The structural claims from Theorem 2.4 remain unchanged:
Therefore,
and hence
Competitive Ratio as a function of d and B
From (5), for all sufficiently large $d$,
$t\;\geq\;\frac{1}{2}\left(\frac{d}{2B\ln d}\right)^{\!\frac{1}{B+1}},$
and thus every online algorithm has competitive ratio at least
$\Omega\!\left(\left(\frac{d}{\ln d}\right)^{\!\frac{1}{B+1}}\right).$
E.2 Refined lower bounds for the unit-capacity case
Here we show how to remove the $\ln m$ loss in this special case by using Poljak’s explicit bound (Theorem 3 in Poljak and Tuza [1989]) and prove the following theorem.
Theorem E.3.
For the single-minded setting with $m$ items and unit capacities, every online algorithm has competitive ratio $\Omega\left(m^{1/3}\right)$.
Poljak’s theorem states that if $t$ is a prime power and $t^{2}-t$ divides ntn-t, then
@@ E.2 Refined lower bounds for the unit-capacity case
We restrict to the case where the number of items is a cube of a prime, $m=n=p^{3}$ for a sufficiently large prime $p$, and set $t\coloneqq p,\ r\coloneqq 2$. Since $t=p$ is a prime (hence a prime power) and $t^{2}-t=p^{2}-p$ divides $n-t=p^{3}-p=p(p^{2}-1)=p(p-1)(p+1)$, the condition of Poljak’s Poljak and Tuza [1989] Theorem 3 holds, and we obtain
$g(n,t,2)\;\geq\;t^{\frac{n-t}{t^{2}-t}}\;=\;p^{\frac{p^{3}-p}{p^{2}-p}}=p^{\frac{p(p^{2}-1)}{p(p-1)}}=p^{p+1}=t^{t+1}.$
@@
We now reuse exactly the same construction and analysis as in the proof of Theorem 2.4, specialized to $B=1$ ($r=2$), with this choice of $m$ and $t$. Lemma F.4 gives $\mathbb{P}\!\left[X\geq 1\right]\;\geq\;0.5$, since $Np=B$.
The structural part is exactly as in the $B$-copies proof, with $B=1$ and $r=2$: the analogue of Claim 10.4 (for $B=1$) says that no feasible allocation can serve buyers from more than one group; the analogue of Claim 10.5 (with $B=1$) says that if $E_{i}$ holds, the offline optimum can serve all $t$ buyers in group $i$. Thus, repeating the offline argument from Theorem 2.4 with $B=1$,
@@
Similarly, the Rubinstein-style online argument from the proof of Theorem 2.4 with $B=1$ shows that each group has total expected value $1$ and, once the algorithm takes value from a group, the expected remaining value in that group is at most 1; combined with the fact that at most one group can be partially served when $B=1$, this yields
@@
which proves Theorem E.3.
Leveraging the recent work of Saxena–Velusamy–Weinberg Saxena et al. [2023] and Alon-Alweiss Alon and Alweiss [2020] we can actually improve slightly the above lower bound.
Theorem E.4.
For the single-minded setting with $m$ items and unit capacities, every online algorithm has competitive ratio $m^{\,1/3+\Omega(1/\log\log m)}$.
We first note that one can obtain a lower bound of this form by translating the lower-bound instance of Saxena–Velusamy–Weinberg Saxena et al. [2023] for prophet inequalities over $q$ partition matroids into a $q$-single-minded instance and counting the number of items.
Instead of going through this translation, we give a direct combinatorial proof, by connecting the Alon–Alweiss measure (AAM) to qualitatively independent partitions and then plugging the resulting partitions into the single-minded lower-bound template used in the proof of Theorem E.3.
@@
Let $\ell:=\textsc{AAM}(p,N)$. By definition there exists a $\mathbb{Z}_{p}$-covering $(\ell,p)$-family $\mathcal{C}\subseteq\mathbb{Z}_{p}^{\ell}$ with $|\mathcal{C}|\geq N$, i.e. for all distinct $u,v\in\mathcal{C}$ and all $s\in\mathbb{Z}_{p}$ there is $i\in[\ell]$ such that $u_{i}-v_{i}\equiv s\pmod{p}$.
Set $X:=[\ell]\times\mathbb{Z}_{p}$, so $|X|=p\ell$. For each $u=(u_{1},\dots,u_{\ell})\in\mathcal{C}$ define a $p$-partition $\mathcal{P}_{u}$ of $X$ by blocks
@@
hence $(i,a+u_{i})=(i,b+v_{i})\in B_{v}(b)$ and so $B_{u}(a)\cap B_{v}(b)\neq\emptyset$. Since $a,b$ were arbitrary, $\mathcal{P}_{u}$ and $\mathcal{P}_{v}$ are qualitatively independent.
Thus $\{\mathcal{P}_{u}:u\in\mathcal{C}\}$ is a family of $|\mathcal{C}|\geq N$ pairwise qualitatively independent $p$-partitions of an $n=p\ell=p\cdot\textsc{AAM}(p,N)$ element set. Therefore $g\bigl(p\cdot\textsc{AAM}(p,N),\,p,\,2\bigr)\geq N$. ∎
Saxena–Velusamy–Weinberg Saxena et al. [2023] show that there exists an absolute constant $\alpha>0$ such that for all sufficiently large primes $p$,
@@
We then instantiate exactly the single-minded lower-bound construction from the proof of Theorem E.3 for $B=1$ using this family $\mathcal{F}$.
The structural and probabilistic analysis of this instance (number of “fully active” groups, offline prophet value, and the Rubinstein-style bound on any online algorithm) is identical to the $B=1$ case of Theorem E.3 and implies:
@@
Appendix F Deferred Proofs
F.1 Fractional Relaxation and Offline Optimum
We prove the following lemma in order to relate the fractional relaxation and the offline optimum.